OpenStudy (anonymous):

a student can eat 4.0 M&Ms every 1.00 seconds. If an M&M has a mass of 63g, determine how many kilograms of M&Ms can be eaten by a class of 20 students in 3.75 hours.

OpenStudy (aaronq):

First find the mass consumed per student per second rate =1 m&m = 63 g rate = 4 m&m/s= 63*4 g/s =252 g/s Now, this rate is in sec but the time interval is in hours, so we need to convert one or the other. \(\sf 3.75~hr*\dfrac{60~min}{1~hr}*\dfrac{60~s}{1 ~min}=13500~s\) The total amount of mass consumed by 1 student in 3.75 hours is rate*time=252 g/s *13500 s = 3402000 g The amount consumed by the class of 20 is 3402000 g * 20=68040000 g = \(6.804*10^{7}~g\)